UFO Company has issued a perpetual bond paying $50 coupon interests each year indefinitely. This perpetual bond, with a face value of $1000, is currently trading at a price of 1,500.
Calculate the current interest rate of UFO’s perpetual bond.
1
Expert's answer
2019-12-03T10:03:53-0500

Current yield=(Annual interest based on coupon rate/Current market price of bond)x100

=1500=50/Current market price of bond)x100

=Current market price of bond=(50/1500)x100

Current market price of bond=3.33%

Thus

The current interest rate of UFO’s perpetual bond=3.33%


Note: Value of bond is decreases from 5% to 3.33%

Initial coupon interest rate=50/1000 x 100=5%

Now its is 3.33%